OpenStudy (anonymous):

ABCD is a rectangle. What is the value of x 28 45 53 98

OpenStudy (camerondoherty):

Do you have an image? c:

OpenStudy (anonymous):

yes going to post it now.

OpenStudy (camerondoherty):

Do you know the pythagorean theorem?

OpenStudy (wil476003):

The answer is 28

OpenStudy (anonymous):

I forgot it. I use to be really good at it.

OpenStudy (wil476003):

You do 53^2=45^2+x^2 solve for x then

OpenStudy (camerondoherty):

Alright, so here's how it goes a^2 + b^2 = c^2 Where c is the hypotneuse and a & b are the sides Note: this can only be used in right triangles So, a^2 is 45^2 we do not have b ^2, it is what we're solving for we have c^2, which is 53^2 So we do this: 53^2 - 45^2 = b^2\ Can you solve that?

OpenStudy (anonymous):

the answer would be 28.

OpenStudy (camerondoherty):

Yep, great job c: